OpenStudy (anonymous):

Lola needs to sign 96 invitations. Using a stopwatch that measures time to tenths of a second, it takes Lola 5.3 seconds to sign her full name. Going by the accuracy of the stopwatch, which is the most accurate determination for the number of minutes Lola needs to sign all 96 invitations? 3.3 minutes 3.3125 minutes 8.48 minutes 8.5 minutes

OpenStudy (kamibug):

Multiply 5.6 by 96. This will give you the total number of seconds it will take her. But, the question wants your answer in minutes. So, after multiplying, divide what you get by 60. :) \(\huge\bf\color{seagreen}{Welcome~to~Openstudy!}\)

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

96*5.3=508.8/60=8.48 mins

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np